### 23 Feb 2024 v1

@@ 4.1 Experimental Settings
Dataset and Evaluation. Following the setup in previous works [34, 62], we conduct experiments on distinct image restoration tasks, including image super-resolution (i.e., classic SR and lightweight SR) and real image denoising. For classic and lightweight image super-resolution, we employ DIV2K [53] and Flickr2K [35] as our training datasets, and use Set5 [3], Set14 [61], B100 [41], Urban100 [23], and Manga109 [42] to evaluate the effectiveness. For real image denoising, we train our model with 320 high-resolution images from SIDD [1] datasets, and use the SIDD test set and DND  [49] dataset for testing. Following [34, 68], we denote the model as MambaIR+ when self-ensemble strategy [35] is used in testing. The performance is evaluated using PSNR and SSIM on the Y channel from the YCbCr color space.
Training Details. In accordance with previous work [34, 7, 62], we perform data augmentation on the training data by applying horizontal flips and random rotations of $90^{\circ},180^{\circ}$, and $270^{\circ}$. Additionally, we crop the original images into $64\times 64$ patches for image SR. For the classic SR, we use the pre-trained weights from the $\times$2 model to initialize those of $\times$3 and $\times$4, and halve the learning rate and total training iterations to reduce training time [35]. To ensure a fair comparison, we adjust the training batch size to 32 for image SR. We employ the Adam [28] as the optimizer for training our MambaIR with $\beta_{1}=0.9,\beta_{2}=0.999$. The initial learning rate is set at $2\times 10^{-4}$ and is halved when the training iteration reaches specific milestones. Our MambaIR model is trained with 8 NVIDIA V100 GPUs.
Table 1: Effects of different design choices of RSSB.
settings Set5 Set14 B100 Urban100 Manga109
(0)baseline 38.57 34.67 32.58 34.15 40.28
(1)+w/o Conv 38.48 34.54 32.56 34.04 40.20
(2)+w/o Conv+CA 38.55 34.64 32.57 34.06 40.14
(3)+replace with MLP 38.55 34.68 32.59 34.22 40.13
@@ 4.2 Ablation Study
Effects of different designs of RSSB. As the core component, the RSSB can improve Mamba with restoration-specific priors. In this section, we ablate different components of the RSSB. The PSNR results of ×$\times 2$ classic SR, presented in Tab. 1, indicate that (1) Exploiting the local patch recurrence in the image can assist the long-range spatial modeling of Mamaba, without which will cause a performance drop. (2) Without using local convolution and channel attention, i.e., directly employing off-the-shelf Mamba for restoration, can only obtain sub-optimal results due to the domain gap. (3) Replace Conv+ChanelAttention with MLP, whose structure will be similar to Transformer, also leads to unfavorable results, indicating that although both SSMs and Attention have the global modeling ability, different assisting modules should be considered for further improvements.
4.3 Image Super-Resolution.
Table 2: Quantitative comparison on classic image super-resolution with state-of-the-art methods.he best and the second best results are in red and blue.
@@
Figure 3: Computational complexity comparison between different methods. We set the standard attention [54] which has a global receptive field as baseline, and denote it as “Full-attn”. We adjust the model to ensure the GPU usage is roughly similar at the beginning, and then scale the input resolution from $48\times 48$ to $84\times 84$.
Figure 4: Qualitative comparison of our MambaIR with CNN- and Transformer-based methods on classic image SR with scale $\times$4.
4.3.1 Classic Image Super-Resolution.
@@
Tab. 2 shows the quantitative results between MambaIR and state-of-the-art super-resolution methods. Thanks to the significant global effective receptive filed, our proposed MambaIR achieves the best performance on almost all five benchmark datasets for all scale factors. For example, our Mamba-based baseline outperforms the Transformer-based benchmark model SwinIR by 0.41dB on Manga109 for ×$\times 2$ scale, demonstrating the prospect of Mamba for image restoration. As for the computational complexity(see Fig. 3), our method is far more efficient than the full-attention baseline [54], and exhibits linear complexity with input resolution which is similar to the efficient attention design such as SwinIR. This observation suggests that out MambaIR has similar scale properties as shifted window attention, and at the same time possesses a global receptive field similar to standard attention. We also give visual comparisons in Fig. 4, and it can be seen that our method can facilitate the reconstruction of sharp edges and natural textures. Moreover, the LAM visualization [20] shown in Fig. 5 demonstrates our MambaIR can utilize more pixels during restoration due to the powerful long-range modeling capability of Mamba.
Figure 5: LAM visualization [20] with CNN and Transformer based methods on $\times$4 SR task. The LAM attribution reflects the importance of each pixel in the input LR image when reconstructing the patch marked with a box. The diffusion index (DI) reflects the range of involved pixels. A higher DI represents a wider range of utilized pixels.
@@
Ours $\times$2 530K 122.1G 38.15 0.9610 33.84 0.9207 32.31 0.9013 32.86 0.9343 39.35 0.9786
@@
Ours $\times$3 538K 55.0G 34.62 0.9286 30.54 0.8459 29.23 0.8083 28.73 0.8635 34.26 0.9482
@@
Ours $\times$4 549K 31.6G 32.41 0.8974 28.76 0.7849 27.70 0.7400 26.53 0.7983 31.05 0.9144
4.3.2 Lightweight Image Super-Resolution.
To demonstrate the scalability of our method, we train the Mamba-light model and compare it with state-of-the-art lightweight image SR methods. Following previous works [71, 39], we also report the number of parameters (#param) and MACs (upscaling a low-resolution image to $1280\times 720$ resolution on all scales). Tab. 3 shows the results. It can be seen that our MambaIR-light outperforms SwinIR-light [34] by up to 0.23dB PSNR on the ×$\times 2$ scale Manga109 dataset with fewer parameters and MACs. The performance results demonstrate the scalability and efficiency of our method.
4.4 Image Denoising
4.4.1 Real Image Denoising.
We further turn to the real image denoising task to evaluate the robustness of our MambaIR when facing real-world degradation. Following [59], we adopt the progressive training strategy for fair comparison. The results are shown in Tab. 4. As one can see, our method achieves comparable performance with existing state-of-the-art models Restormer [57] and outperforms other methods such as Uformer [57] by 0.12dB PSNR on SIDD dataset. The promising results indicate the ability of our method in real image denoising.
@@ 4.4 Image Denoising
5 Limitation and Future Work
While MambaIR appears as a competitive restoration backbone against mainstream Transformer-based methods, it can be further improved with restoration-specific designs. For example, in the proposed method, the local similarity and channel interaction priors are employed to obtain better feature representation. However, other restoration prior such as high-frequency enhancement can be also used. As a pure SSM-based method, our MambaIR does not include the Attention module, therefore how to appropriately combine Attention and Mamba remains an interesting research topic. Moreover, although this work has covered multiple image restoration tasks, some other tasks such as image deblurring and deraining can also be explored in the future. Finally, despite the promising results shown, we would like to point out that the Mamba-based image restoration network is still in its early stages. With the increasing interest in Mamba, it will be promising to study the state-space models for low-level vision.
@@ 6 Conclusion
In this work, we explore for the first time the power of the recent advanced state space model i.e., Mamba for image restoration through extensive experiments. By exploiting restoration-specific prior, we propose a brand-new backbone solution, named MambaIR, that can well capture the local patch recurrence and channel interaction while maintaining strong long-range capability and linear complexity, serving as a simple but effective Mamba-based benchmark model for image restoration. Extensive experiments on multiple restoration tasks demonstrate the effectiveness of MambaIR against existing CNN- and Tansformer-based methods.

### 25 Mar 2024 v2

@@ 4.1 Experimental Settings
Dataset and Evaluation. Following the setup in previous works [36, 66], we conduct experiments on various image restoration tasks, including image super-resolution (i.e., classic SR, lightweight SR, and real SR) and image denoising (i.e., Gaussian color image denoising and real-world denoising). We employ DIV2K [56] and Flickr2K [37] to train classic SR models and use only DIV2K to train lightweight SR models. Moreover, we use Set5 [4], Set14 [65], B100 [44], Urban100 [24], and Manga109 [45] to evaluate the effectiveness of different SR methods. For gaussian color image denoising, we utilize DIV2K [56], Flickr2K [37], BSD500 [3], and WED [43] as our training datasets. Our testing datasets for guassian color image denoising includes BSD68 [44], Kodak24 [16], McMaster [74], and Urban100 [24]. For real image denoising, we train our model with 320 high-resolution images from SIDD [1] datasets, and use the SIDD test set and DND  [52] dataset for testing. Following [36, 76], we denote the model as MambaIR+ when self-ensemble strategy [37] is used in testing. The performance is evaluated using PSNR and SSIM on the Y channel from the YCbCr color space. In addition, we also include other restoration tasks such as JPEG compression artifact reduction, more experimental results can be seen in the supplementary material.
Table 1: Ablation experiments for different design choices of RSSB.
settings #param MACs Set5 Set14 B100 Urban100 Manga109
(0)baseline 16.7M 439G 38.57 34.67 32.58 34.15 40.28
(1)+w/o Conv 11.9M 316G 38.48 34.54 32.56 34.04 40.20
(2)+w/o Conv+CA 11.8M 315G 38.55 34.64 32.57 34.06 40.14
(3)+replace with MLP 14.3M 379G 38.55 34.68 32.59 34.22 40.13
Training Details. In accordance with previous works [36, 9, 66], we perform data augmentation by applying horizontal flips and random rotations of $90^{\circ},180^{\circ}$, and $270^{\circ}$. Additionally, we crop the original images into $64\times 64$ patches for image SR and $128\times 128$ patches for image denoising during training. For image SR, we use the pre-trained weights from the $\times$2 model to initialize those of $\times$3 and $\times$4 and halve the learning rate and total training iterations to reduce training time [37]. To ensure a fair comparison, we adjust the training batch size to 32 for image SR and 16 for image denoising. We employ the Adam [30] as the optimizer for training our MambaIR with $\beta_{1}=0.9,\beta_{2}=0.999$. The initial learning rate is set at $2\times 10^{-4}$ and is halved when the training iteration reaches specific milestones. Our MambaIR model is trained with 8 NVIDIA V100 GPUs.
@@ 4.2 Ablation Study
Effects of different designs of RSSB. As the core component, the RSSB can improve Mamba with restoration-specific priors. In this section, we ablate different components of the RSSB. The results, presented in Tab. 1, indicate that (1) directly using the standard Mamba to process flattened sequences can lead to local pixel forgetting, and the utilization of simple convolution layers can effectively enhance the local interaction. (2) Without using additional convolution and channel attention, i.e., directly employing off-the-shelf Mamba for restoration, can only obtain sub-optimal results, which also supports our previous analysis. (3) Replacing Conv+ChanelAttention with MLP, whose resulted structure will be similar to Transformer, also leads to unfavorable results, indicating that although both SSMs and Attention have the global modeling ability, the behavior of these two modules is different and thus accustomed block structure should be considered for further improvements.
Figure 4: Qualitative comparison of our MambaIR with CNN and Transformer based methods on classic image SR with scale $\times$4.
Table 2: Ablation experiments for different scan modes in VSSM.
scan mode #param MACs Set5 Set14 B100 Urban100 Manga109
one-direction 16.7M 439G 38.53 34.63 32.58 34.06 40.31
two-direction 16.7M 439G 38.56 34.60 32.56 33.96 40.14
baseline 16.7M 439G 38.57 34.67 32.58 34.15 40.28
Table 3: Quantitative comparison on classic image super-resolution with state-of-the-art methods. The best and the second best results are in red and blue.
@@
Effects of Different Scan Modes in VSSM. To allow Mamba to process 2D images, the feature map needs to be flattened before being iterated by the state-space equation. Therefore, the unfolding strategy is particularly important. In this work, we follow  [38] which uses scans in four different directions to generate scanned sequences. Here, we ablate different scan modes to study the effects, the results are shown in Tab. 2. Compared with one-direction(top-left to bottom-right) and two-direction(top-left to bottom-right, bottom-right to top-left), using four directions of scanning allows the anchor pixel to perceive a wider range of neighborhoods, thus achieving better results. Surprisingly, using more scanning directions brings an almost negligible increase in the number of parameters and computational complexity. This benefit may facilitate future exploration of more effective scanning mechanisms for the Mamba-based image restoration networks. We also include other ablation experiments, such as the layer number of RSSBs, please see supplementary material for more analysis.
4.3 Comparison on Image Super-Resolution
@@
Classic Image Super-Resolution. Tab. 3 shows the quantitative results between MambaIR and state-of-the-art super-resolution methods. Thanks to the significant global receptive field, our proposed MambaIR achieves the best performance on almost all five benchmark datasets for all scale factors. For example, our Mamba-based baseline outperforms the Transformer-based benchmark model SwinIR by 0.41dB on Manga109 for ×$\times 2$ scale, demonstrating the prospect of Mamba for image restoration. We also give visual comparisons in Fig. 4, and it can be seen that our method can facilitate the reconstruction of sharp edges and natural textures.
Table 4: Model size comparisons. We compare the complexity on the ×$\times 4$ classic SR model of different methods, the output size is set to $3\times 640\times 640$ for MACs.
Method EDSR [37] RCAN  [76] SRFBN  [35] HAN  [51] CSNLN  [48] SwinIR  [36] MambaIR (Ours)
Params (M) 43.1 15.6 3.6 16.1 7.2 11.9 16.7
Mult-Adds $(\mathrm{G})$ 1,286 407 498 420 103,640 336 439
PSNR on Urban100 (dB) 26.64 26.82 26.60 26.85 27.22 27.45 27.68
PSNR on Manga109 (dB) 31.02 31.22 31.15 31.42 31.43 32.03 32.32
Figure 5: Computational complexity comparison with different input scales. We set the standard attention [57] which has a global receptive field as baseline, and denote it as "Full-attn". We adjust the model to ensure the GPU usage is roughly similar at the beginning, and then scale the input resolution from $48\times 48$ to $84\times 84$.
Model Complexity Comparison.  Tab. 4 provides the comparison of different methods on parameters and MACs. Our MambaIR has comparable parameters and MACs compared with SwinIR, but can achieve a global receptive field and better PSNR performance. We further give comparison results of computational complexity with input sizes in  Fig. 5. As one can see, our method is far more efficient than the full-attention baseline [57] and exhibits linear complexity with input resolution which is similar to the efficient attention techniques such as SwinIR. These observations above suggest that out MambaIR has similar scale properties as shifted window attention, while possessing a global receptive field similar to standard full attention.
Lightweight Image Super-Resolution. To demonstrate the scalability of our method, we train the Mamba-light model and compare it with state-of-the-art lightweight image SR methods. Following previous works [80, 41], we also report the number of parameters (#param) and MACs (upscaling a low-resolution image to $1280\times 720$ resolution). Tab. 5 shows the results. It can be seen that our MambaIR-light outperforms SwinIR-light [36] by up to 0.17dB PSNR on the ×$\times 4$ scale Manga109 dataset with similar parameters and MACs. The performance results demonstrate the scalability and efficiency of our method.
@@
SRFormer-light [80] $\times$2 853K 236G 38.23 0.9613 33.94 0.9209 32.36 0.9019 32.91 0.9353 39.28 0.9785
Ours $\times$2 859K 198.1G 38.16 0.9610 34.00 0.9212 32.34 0.9017 32.92 0.9356 39.31 0.9779
@@
SRFormer-light [80] $\times$3 861K 105G 34.67 0.9296 30.57 0.8469 29.26 0.8099 28.81 0.8655 34.19 0.9489
Ours $\times$3 867K 88.7G 34.72 0.9296 30.63 0.8475 29.29 0.8099 29.00 0.8689 34.39 0.9495
@@
SRFormer-light [80] $\times$4 873K 62.8G 32.51 0.8988 28.82 0.7872 27.73 0.7422 26.67 0.8032 31.17 0.9165
Ours $\times$4 879K 50.6G 32.51 0.8993 28.85 0.7876 27.75 0.7423 26.75 0.8051 31.26 0.9175
Real-world Image Super-resolution. We also investigate the performance of the network for real-world image restoration. We follow the training protocol in  [9] to train our MambaIR-real model. Since there are no ground-truth images for this task, only the visual comparison is given in Fig. 6. Compared with the other methods, our MambaIR exhibits a notable advancement in resolving fine details and texture preservation, demonstrating the robustness of our method.
Figure 6: Qualitative comparison with RealSR [27], ESRGAN [60], BSRGAN [69], Real-ESRGAN [59], and SwinIR [36] on real image super-resolution with scale $\times$4.
4.4 Comparison on Image Denoising
Gaussian Color Image Denoising. The results of gaussian color image denoising are shown in  Tab. 6. Following [67, 70], the compared noise levels include 15, 25 and 50. As one can see, our model achieves better performance than all compared methods. In particular, it surpasses the SwinIR [36] by up to 0.25dB with $\sigma$=50 on the Urban100 dataset. We also give a visual comparison in Fig. 7. Thanks to the global receptive field, our MambaIR can achieve better structure preservation, leading to clearer edges and natural shapes.
Table 6: Quantitative comparison on gaussian color image denoising with state-of-the-art methods.
Method BSD68 Kodak24 McMaster Urban100
$\sigma$=15 $\sigma$=25 $\sigma$=50 $\sigma$=15 $\sigma$=25 $\sigma$=50 $\sigma$=15 $\sigma$=25 $\sigma$=50 $\sigma$=15 $\sigma$=25 $\sigma$=50
IRCNN [72] 33.86 31.16 27.86 34.69 32.18 28.93 34.58 32.18 28.91 33.78 31.20 27.70
FFDNet [73] 33.87 31.21 27.96 34.63 32.13 28.98 34.66 32.35 29.18 33.83 31.40 28.05
DnCNN [71] 33.90 31.24 27.95 34.60 32.14 28.95 33.45 31.52 28.62 32.98 30.81 27.59
DRUNet [68] 34.30 31.69 28.51 35.31 32.89 29.86 35.40 33.14 30.08 34.81 32.60 29.61
SwinIR [36] 34.42 31.78 28.56 35.34 32.89 29.79 35.61 33.20 30.22 35.13 32.90 29.82
Restormer [63] 34.40 31.79 28.60 35.47 33.04 30.01 35.61 33.34 30.30 35.13 32.96 30.02
MambaIR 34.43 31.80 28.61 35.34 32.91 29.85 35.62 33.35 30.31 35.17 32.99 30.07
MambaIR+ 34.44 31.82 28.64 35.35 32.92 29.87 35.63 33.36 30.32 35.17 32.99 30.06
Figure 7: Qualitative comparison of our MambaIR with other methods on color image denoising task with noise level level $\sigma$=50.
Real Image Denoising. We further turn to the real image denoising task to evaluate the robustness of MambaIR when facing real-world degradation. Following [63], we adopt the progressive training strategy for fair comparison. The results, shown in Tab. 7, suggest that our method achieves comparable performance with existing state-of-the-art models Restormer [61] and outperforms other methods such as Uformer [61] by 0.12dB PSNR on SIDD dataset. The promising results indicate the ability of our method in real image denoising.
@@ 5 Conclusion
In this work, we explore for the first time the power of the recent advanced state space model, i.e., Mamba, for image restoration, to help resolve the dilemma of trade-off between efficient computation and global effective receptive field. Specifically, we introduce the local enhancement to mitigate the neighborhood pixel forgetting problem from the flattening strategy and propose channel attention to reduce channel redundancy. Extensive experiments on multiple restoration tasks demonstrate our MambaIR serves as a simple but effective state-space model for image restoration.